# Generalized Pinching-Antenna Systems: A Tutorial on Principles, Design Strategies, and Future Directions

### v1

@@ III-C Uplink Pinching-Antenna Systems
Multiple SDMA Antenna position, beamforming Penalty method + SCA [67, 68]; deep learning [69, 70]
@@ IV-A Enabling ISAC with Pinching Antennas
ISAC has emerged as a transformative technology for next-generation wireless networks, enabling seamless integration of communication and sensing functionalities in a unified infrastructure [89, 90, 91, 92]. By jointly utilizing spectrum, hardware, and signal processing resources, ISAC is expected to support a wide range of applications, such as high-precision localization, environment monitoring, and smart industrial automation. Recent research has identified ISAC as a key use case for 6G and beyond, attracting significant attention from both academia and industry [1].
@@
$0\leq\tilde{x}_{T,p}\leq\tilde{x}_{\max}.\forall p,$ (31d)
@@ IV-B Facilitating Cooperative Communications with Pinching Antennas
Instead of solving problem (34) directly, the authors of [107] instead explored the characteristics of cooperative transmission to gain insights that can facilitate its solution. Adopting different levels of cooperation, the joint BS and the pinching-antenna transmission system can operate in three district modes: 1) standalone deployment (SD): the BS distributes its signal to multiple waveguide processing units (WPUs), each of which are deployed at a separate location. Each WPU independently feeds the signal to pinching antennas along its associated waveguide, thereby enhancing the system coverage and communication performance; 2) semi-cooperative deployment (SCD): the BS transmits its signal to a centralized WPU located away from the BS. This WPU then simultaneously distributes the signal to multiple waveguides, enabling coordinated transmission with moderate complexity; 3) fully-cooperative deployment (FCD): the WPU is fully integrated in the BS, allowing the signal to be fed directly into the waveguides without additional intermediate interfaces. This tightly coupled configuration offers the highest level of integration and potentially the lowest transmission latency.
Along with the three cooperative transmission modes in [107], power allocation in the SD and SCD schemes is performed in proportion to the number of RF chains at the BS and the waveguides for meeting the total transmit power budget across different transmission schemes. In the FCD scheme, the BS and pinching antennas perform joint signal processing and can flexibly allocate power among the $N_{B}+P$ RF chains based on the full channel state information (CSI) of both the channels from the BS to the users and the pinching antennas to the user. Following the pinching antenna placement scheme derived in [52], the received SNR for different cooperative transmission schemes can be obtained in closed-form as shown in Table VI, where $L_{G}$ is the distance between the pinching antenna and the user, $\xi_{P}$ is the path loss coefficient, and $\tilde{\gamma}=\frac{P_{\max}}{\sigma_{s}^{2}}$ denotes the transmit SNR. The received SNR for joint transmission is influenced by multiple factors, and performance gains are not guaranteed under all parameter configurations. Hence, the choice of an appropriate cooperation scheme has to be tailored to the specific communication context.
TABLE VI: Received SNR for different schemes [107]
@@
It is worth noting that deploying pinching antennas in existing cellular networks inherently involves multi-point cooperative transmission and support for multiple users, which leads to significantly more complex optimization problems. These include joint beamforming design, power allocation, scheduling, and user association, all of which collectively increase the problem’s dimensionality and computational complexity. Deployment strategies aimed at maximizing the performance gains of multi-user pinching-antenna systems have been studied in [36, 80, 66], and can be integrated with coordinated multipoint systems [108, 109] to achieve further performance improvements. Therefore, effective optimization is critical for fully realizing the benefits of cooperative transmission in practical network deployments, particularly with regard to spectral efficiency, coverage enhancement, and scalability.
@@ IV-C1 Pinching Antennas-Assisted Physical Layer Security
By leveraging the unique capability of dynamically configuring antenna positions along dielectric waveguides, pinching-antenna systems have recently emerged as a promising paradigm for enhancing physical layer security in wireless networks. The spatial flexibility introduces new degrees of freedom for secure transmission, allowing transmitters to form strong LoS links to legitimate users while spatially distancing or nulling eavesdroppers. Recent studies have explored a range of secure transmission strategies for pinching-antenna systems [110, 111, 112, 113, 114, 115, 116]. For instance, the work [110] formulated joint baseband and pinching beamforming optimization problems to maximize the secrecy rate and proposed gradient-based and FP-BCD algorithms, showing that the pinching-antenna systems outperform conventional fixed-antenna systems as well as advanced linear precoding benchmarks in both single-user and multi-user wiretap scenarios. On the other hand, a rigorous performance analysis for single-user systems was conducted in [111], where closed-form expressions for the average secrecy capacity and secrecy outage probability were derived. Besides, the results also demonstrated that placing pinching antennas close to legitimate users and increasing spatial separation from eavesdroppers significantly enhances secrecy metrics. Other works have further extended the study to artificial noise injection [112], and secure multi-user MIMO scenarios [113]. In [115], the authors investigated physical-layer security in dual-waveguide pinching-antenna systems, where multiple pinching antennas act as flexible channel reconfigurators to enhance secure sum rate and energy efficiency, and proposes a two-stage optimization algorithm that jointly designs antenna placement, beamforming, and artificial noise under parallel and orthogonal waveguide deployments. In addition, the authors of [116] leveraged the spatial reconfigurability of pinching-antenna systems to boost covert communication capabilities, demonstrating that dynamic antenna positioning and beamforming can maximize covert rates and enhance transmission stealthiness in the presence of adversaries. Overall, the existing research results show that pinching-antenna-enabled flexible antenna placement, joint secure beamforming, and waveguide-based LoS links together provide powerful tools for improving secrecy performance in future wireless systems, though challenges remain in real-time optimization, practical hardware design, and the integration of security with other advanced wireless functionalities.
@@ IV-C2 Pinching Antennas for Wireless Power Transfer
Pinching-antenna systems have recently gained attention as a powerful tool for overcoming critical limitations in wireless power transfer (WPT) and energy harvesting networks. Traditional WPT systems often suffer from the “double near-far” problem, where users far away from the transmitter receive both less energy and have weaker communication channels. By enabling the flexible activation and placement of antennas along waveguides, pinching-antenna systems allow transmitters to deliver energy much closer to distant users, significantly improving energy harvesting efficiency and uplink throughput. For example, the work [117] proposed a wireless powered pinching-antenna network, where distributed pinching antennas are jointly optimized for downlink energy transfer and uplink data collection. This architecture enhances user fairness and network performance compared to conventional fixed-antenna solutions. Building on these ideas, further works have investigated resource allocation, optimal antenna activation, and joint power optimization in WPT-enabled pinching-antenna systems, using advanced methods such as game theory and particle swarm optimization to efficiently handle the associated non-convex problems [118, 119]. The wireless powered mobile edge computing in pinching-antenna system was investigated in [120].
@@ IV-C3 More Applications of Pinching Antennas
Recent advances have also demonstrated that pinching-antenna systems hold significant promise for a wide range of emerging wireless applications beyond classical communications. For example, the authors of [125] introduced a novel pinching-antenna-assisted index modulation scheme, where both conventional modulation symbols and the indices of activated pinching-antenna positions are used for information delivery. This approach enables higher spectral efficiency without increasing hardware complexity, and substantially outperforms conventional modulation systems, highlighting the versatility and potential of pinching antennas in enabling advanced modulation and coding techniques. Pinching antenna-assisted symbiotic radio was studied in [126], aiming to enhance both primary and secondary transmissions by leveraging the flexibility of pinching antennas. While, the authors of [127] investigated the pinching-antenna enabled indoor positioning.
@@ V Charting the Future: Open Questions and Research Directions
V-A Pinching-Antenna System Design under LoS Blockage
LoS blockage is a major challenge in high-frequency wireless systems, particularly in complex indoor environments, dense urban areas, and tunnel-like structures. Obstacles and user mobility can obstruct the direct propagation path between transmitters and receivers, leading to significant signal attenuation and degraded communication performance. These effects are especially pronounced in systems that rely on direct-path transmission. As pinching-antenna systems are envisioned for deployment in such environments, it is crucial to understand, mitigate, and even exploit the impact of LoS blockage for robust system design.

### v2

@@ III-B2 Multiple Pinching Antennas Per Waveguide
Multiple SDMA Antenna position, beamforming Penalty method + SCA [73, 74]; deep learning [75, 76, 93]
III-C Pinching-Antenna System Design for Random LoS and NLoS Channels
In the previous subsections, we considered downlink pinching-antenna system design assuming the availability of an LoS channel. However, in more complex wireless environments, the LoS channel can be intermittently blocked by obstacles, and the NLoS component can dominate due to multipath scattering. Therefore, it is important to study the performance of pinching-antenna systems for random LoS blockage and NLoS fading, as these factors significantly affect system performance, especially in dynamic environments such as urban canyons or indoor areas with heavy obstruction. In this subsection, we investigate pinching-antenna system design for random LoS and NLoS channels.
III-C1 System Model and Channel Setup
Consider a basic downlink system where a single pinching antenna serves multiple users using TDMA. The users are randomly deployed in a rectangular area of size $D_{x}\times D_{y}$. In TDMA, users are allocated orthogonal time slots, ensuring that interference is minimized. The position of the pinching antenna is optimized once for the whole scheduling duration for all users. Without loss of generality, we consider a communication region where $M$ users are randomly distributed.
The channel between the pinching antenna and user $m$ is modeled as a combination of LoS and NLoS components. The total channel gain is given by
$h_{m}=\xi_{m}h_{m}^{\mathrm{LoS}}+h_{m}^{\mathrm{NLoS}},$ (28)
[…]
III-C2 Optimization Problems
To optimize the performance of the pinching-antenna system under random LoS and NLoS channels, we formulate two complementary optimization problems based on different performance metrics:
2.1 Average SNR-based Design: The average-SNR-based design focuses on maximizing the long-term average received SNR of the system by optimizing the antenna placement. We use the average SNR as a tractable surrogate for long-term throughput, since for a fixed modulation and coding scheme, the achievable rate is a monotonically increasing function of the received SNR, and averaging the SNR yields much simpler expressions than averaging $\log(1+\mathrm{SNR})$, which facilitates analysis and optimization. The goal is to maximize the minimum average SNR across all users, so that the system can simultaneously achieve good spectral-efficiency performance and preserve user fairness. The corresponding optimization problem is formulated as
[…]
${\rm subject~to}\quad$ $0\leq\tilde{x}\leq D_{x},$ (29b)
[…]
2.2 Outage-Constrained Design: The outage-constrained design aims to guarantee a minimum level of reliability for each user by constraining the outage probability. Instead of maximizing the instantaneous rate directly, we optimize an SNR threshold that serves as a design target for selecting modulation and coding schemes: a higher threshold allows more aggressive coding and higher rates, while still respecting reliability requirements. The goal is therefore to maximize this SNR threshold while ensuring the probability that the instantaneous SNR falls below the threshold is less than a prescribed value for each user. The outage-constrained problem is formulated as
$\underset{\tilde{x},t}{{\rm maximize}}\quad$ $t$ (30a)
${\rm subject~to}\quad$ $\mathbb{P}\left(\Gamma_{m}(\tilde{x})\geq t\right)\geq 1-\varepsilon_{m},\quad\forall m\in\mathcal{M}$ (30b)
$0\leq\tilde{x}\leq D_{x},$ (30c)
where $t$ is the target SNR threshold, and $\varepsilon_{m}$ is the prescribed outage probability for user $m$.
These two optimization problems capture the trade-off between maximizing throughput (average SNR) and ensuring reliability (outage-constrained SNR) in a stochastic LoS/NLoS environment, where random blockage and fading introduce additional challenges to system design. As shown in [94], problems (29) and (30) can be efficiently solved by using bisection-based algorithms.
III-C3 Simulation Results
Figure 11: CDF of the max-min average SNR for pinching- and fixed-antenna systems with $D_{x}=40$ m.
Figure 12: Optimal SNR threshold $t^{*}$ versus outage probability $\varepsilon_{m}=\varepsilon,\forall m,$ for pinching and fixed antenna systems with $D_{x}=40$ m.
To demonstrate the performance of the pinching-antenna system under random LoS and NLoS conditions with the max-min average SNR metric, we simulate the cumulative distribution functions (CDFs) of the max-min average SNR for both pinching-antenna and fixed-antenna systems, as shown in Fig. 11. In the simulation, we consider different numbers of users, $M\in\{2,4,8\}$ and we set $\mu_{m}^{2}=-80$ dBm, $\sigma_{m}^{2}=-90$ dBm, communication area side length of $D_{y}=10$ m and $\beta=0.001$. As observed, the proposed pinching-antenna system consistently achieves higher SNR across all percentiles compared to its fixed counterpart, indicating a clear improvement in both typical and worst-case performance.
To evaluate the performance of the pinching-antenna system under the outage-constrained metric, we plot the optimal SNR thresholds $t^{*}$ for both pinching- and fixed-antenna systems as a function of outage probability, considering two different LoS blockage coefficients, $\beta$ in Fig. 12. A larger $\beta$ indicates a larger LoS blockage probability. As expected, $t^{*}$ increases monotonically with the outage probability $\varepsilon$ for both systems, as a more relaxed reliability constraint allows higher transmission thresholds. At a given reliability level, the pinching-antenna system consistently outperforms the fixed-antenna system, demonstrating its ability to adaptively reposition and improve the LoS condition for users.
Overall, these results provide quantitative evidence that pinching antennas can significantly enhance link robustness and reliability, even in challenging blockage-dominated environments.
III-D Uplink Pinching-Antenna Systems
Recent literature has begun to address these challenges from several perspectives [59, 56, 85, 95]. The first detailed study of the multi-user uplink in pinching-antenna systems was in [59], where TDMA is used to avoid inter-channel interference. The authors formulated and solved a joint optimization problem to determine the radio resources allocation and pinching-antenna positions to maximize the minimum data rate among users, thereby ensuring fairness and meeting diverse QoS requirements. To further improve spectral efficiency, NOMA-based uplink pinching-antenna systems were studied in [56]. Considering a MIMO setting, the authors of [85] investigated the joint design of both uplink and downlink transmission for pinching-antenna systems. The sum-rate maximization problem was formulated and solved using an iterative algorithm. Simulation results demonstrated the performance gain of pinching-antenna systems in uplink MIMO scenarios as compared to conventional fixed-antenna systems.
To address this issue, a recent study in [95] proposed a segmented waveguide-enabled pinching-antenna architecture. In this design, the traditional long dielectric waveguide is divided into multiple shorter segments, each equipped with its own feed point and associated pinching antennas. By ensuring that only one pinching antenna is active per segment, this approach effectively eliminates inter-antenna radiation leakage and enables a tractable, physically consistent uplink signal model that supports multi-antenna operation. The work further introduced three practical operating protocols, including the segment selection, the segment aggregation, and the segment multiplexing, each offering a distinct trade-off in terms of performance and implementation complexity.
From a practical standpoint, waveguide segmentation also introduces non-negligible implementation implications. A continuous dielectric waveguide represents the lowest-complexity baseline, where a single low-loss dielectric waveguide is deployed along the service area and pinching antennas are realized as local radiating points without additional active components. By contrast, segmented-waveguide solutions require extra RF components, such as connectors, switches, or reconfigurable couplers, to isolate or selectively feed different segments. While this segmentation helps mitigate uplink inter-antenna radiation leakage, it inevitably increases hardware and installation complexity. Moreover, segmentation introduces an additional algorithmic design dimension: The system must decide which waveguide or segment to activate based on the spatial distribution and traffic demands of the users, motivating further research on user-aware waveguide selection and scheduling strategies in practical deployments.
III-E Control and Implementation of Pinching-Antenna Positioning
In the previous subsections, we mainly focused on where the pinching antenna should be placed by solving optimization problems. In practice, these designs must be realized through a control mechanism that adjusts the pinching-antenna position along the waveguide. Depending on the implementation, there are two relevant approaches for implementing pinching-antenna control: continuous-position and discrete-position.
In a continuous-position implementation, the pinching antenna is mounted on a rail or slider, enabling its position along the waveguide to be continuously adjusted. This supports both a real-time user-centric mode of operation and a long-term environment-aware mode of operation. In the real-time mode, the antenna position is adapted according to the instantaneous or short-term estimated locations of the active users, which can be obtained using methods such as cellular positioning, device-based sensing, or machine learning-based localization techniques. In the long-term mode, the antenna position is optimized based on the environment and network-level performance, for example by accounting for the spatial user distribution, blockage patterns, and QoS requirements in the service area. In this case, the antenna position is updated only occasionally when the environment changes, and serves as a slowly varying configuration knob that complements conventional beamforming and resource allocation.
For a discrete-position implementation, multiple pinching antennas are pre-deployed at specific locations along the waveguide. The control task is to decide which antenna(s) to activate based on the real-time user distribution (real-time mode) or based on static environmental factors such as traffic patterns and persistent blockage (long-term mode). This activation-based control can be implemented electronically and supports moderate to fast reconfiguration.
The methods presented in this work primarily correspond to the real-time mode of operation for both continuous and discrete positioning, where antenna positions (or active antennas) are updated dynamically based on user locations or large-scale channel information. In practice, such real-time operation can be realized by segment switching in discrete-position implementations and by small mechanical motions in continuous-position implementations, scheduled over a finite set of candidate activation points and operating on the time scales of typical human or vehicular mobility rather than at the symbol level. At the same time, the same optimization framework can be applied in an environment-aware manner by replacing instantaneous user locations with statistical information (e.g., traffic and blockage patterns), which is particularly attractive for outdoor or highly dynamic scenarios where per-user tracking is difficult and pinching antennas are used to enhance coverage and throughput at the network level rather than to follow individual users.
@@ IV-A Enabling ISAC with Pinching Antennas
ISAC has emerged as a transformative technology for next-generation wireless networks, enabling seamless integration of communication and sensing functionalities in a unified infrastructure [96, 97, 98, 99, 100, 101]. By jointly utilizing spectrum, hardware, and signal processing resources, ISAC is expected to support a wide range of applications, such as high-precision localization, environment monitoring, and smart industrial automation. Recent research has identified ISAC as a key use case for 6G and beyond, attracting significant attention from both academia and industry [1].
@@
$0\leq\tilde{x}_{T,p}\leq\tilde{x}_{\max},\,\forall p,$ (34d)
@@ IV-B Facilitating Cooperative Communications with Pinching Antennas
Instead of solving problem (37) directly, the authors of [116] instead explored the characteristics of cooperative transmission to gain insights that can facilitate its solution. Adopting different levels of cooperation, the joint BS and the pinching-antenna transmission system can operate in three district modes: 1) standalone deployment (SD): the BS distributes its signal to multiple waveguide processing units (WPUs), each of which is deployed at a separate location. Each WPU independently feeds the signal to pinching antennas along its associated waveguide, thereby enhancing the system coverage and communication performance; 2) semi-cooperative deployment (SCD): the BS transmits its signal to a centralized WPU located away from the BS. This WPU then simultaneously distributes the signal to multiple waveguides, enabling coordinated transmission with moderate complexity; 3) fully-cooperative deployment (FCD): the WPU is fully integrated in the BS, allowing the signal to be fed directly into the waveguides without additional intermediate interfaces. This tightly coupled configuration offers the highest level of integration and potentially the lowest transmission latency.
Along with the three cooperative transmission modes in [116], power allocation in the SD and SCD schemes is performed in proportion to the number of RF chains at the BS and the waveguides for meeting the total transmit power budget across different transmission schemes. In the FCD scheme, the BS and pinching antennas perform joint signal processing and can flexibly allocate power among the $N_{B}+P$ RF chains based on the full channel state information (CSI) of both the channels from the BS to the users and the pinching antennas to the user. Following the pinching antenna placement scheme derived in [58], the received SNR for different cooperative transmission schemes can be obtained in closed-form as shown in Table VIII, where $L_{G}$ is the distance between the pinching antenna and the user, $\xi_{P}$ is the path loss coefficient, and $\tilde{\gamma}=\frac{P_{\max}}{\sigma_{s}^{2}}$ denotes the transmit SNR. The received SNR for joint transmission is influenced by multiple factors, and performance gains are not guaranteed under all parameter configurations. Hence, the choice of an appropriate cooperation scheme has to be tailored to the specific communication context.
TABLE VIII: Received SNR for different schemes [116]
@@
It is worth noting that deploying pinching antennas in existing cellular networks inherently involves multipoint cooperative transmission and support for multiple users, which leads to significantly more complex optimization problems. These include joint beamforming design, power allocation, scheduling, and user association, all of which collectively increase the problem’s dimensionality and computational complexity. Deployment strategies aimed at maximizing the performance gains of multi-user pinching-antenna systems have been studied in [41, 86, 72], and can be integrated with coordinated multipoint systems [117, 118] to achieve further performance improvements. Therefore, effective optimization is critical for fully realizing the benefits of cooperative transmission in practical network deployments, particularly with regard to spectral efficiency, coverage enhancement, and scalability.
@@ IV-C1 Pinching Antennas-Assisted Physical Layer Security
By leveraging the unique capability of dynamically configuring antenna positions along dielectric waveguides, pinching-antenna systems have recently emerged as a promising paradigm for enhancing physical layer security in wireless networks. The spatial flexibility introduces new degrees of freedom for secure transmission, allowing transmitters to form strong LoS links to legitimate users while spatially distancing or nulling eavesdroppers. Recent studies have explored a range of secure transmission strategies for pinching-antenna systems [119, 120, 121, 122, 123, 124, 125]. For instance, the work [119] formulated joint baseband and pinching beamforming optimization problems to maximize the secrecy rate and proposed gradient-based and FP-BCD algorithms, showing that the pinching-antenna systems outperform conventional fixed-antenna systems as well as advanced linear precoding benchmarks in both single-user and multi-user wiretap scenarios. On the other hand, a rigorous performance analysis for single-user systems was conducted in [120], where closed-form expressions for the average secrecy capacity and secrecy outage probability were derived. Besides, the results also demonstrated that placing pinching antennas close to legitimate users and increasing spatial separation from eavesdroppers significantly enhances secrecy metrics. Other works have further extended the study to artificial noise injection [121], and secure multi-user MIMO scenarios [122]. In [124], the authors investigated physical-layer security in dual-waveguide pinching-antenna systems, where multiple pinching antennas act as flexible channel reconfigurators to enhance secure sum rate and energy efficiency, and proposed a two-stage optimization algorithm that jointly designs antenna placement, beamforming, and artificial noise under parallel and orthogonal waveguide deployments. In addition, the pinching-antenna-assisted covert communications have been studied in [125, 126]. In particular, the authors of [125] leveraged the spatial reconfigurability of pinching-antenna systems to boost covert communication capabilities, demonstrating that dynamic antenna positioning and beamforming can maximize covert rates and enhance transmission stealthiness in the presence of adversaries. While the authors of [126] studied the pinching-antenna-enabled covert communications in ISAC systems. Overall, the existing research results show that pinching-antenna-enabled flexible antenna placement, joint secure beamforming, and waveguide-based LoS links together provide powerful tools for improving secrecy performance in future wireless systems, though challenges remain in real-time optimization, practical hardware design, and the integration of security with other advanced wireless functionalities.
@@ IV-C2 Pinching Antennas for Wireless Power Transfer
Pinching-antenna systems have recently gained attention as a powerful tool for overcoming critical limitations in wireless power transfer (WPT) and energy harvesting networks. Traditional WPT systems often suffer from the “double near-far” problem, where users far away from the transmitter receive both less energy and have weaker communication channels. By enabling the flexible activation and placement of antennas along waveguides, pinching-antenna systems allow transmitters to deliver energy much closer to distant users, significantly improving energy harvesting efficiency and uplink throughput. For example, the work in[127] proposed a wireless powered pinching-antenna network, where distributed pinching antennas are jointly optimized for downlink energy transfer and uplink data collection. This architecture enhances user fairness and network performance compared to conventional fixed-antenna solutions. Building on these ideas, further works have investigated resource allocation, optimal antenna activation, and joint power optimization in WPT-enabled pinching-antenna systems, using methods such as game theory and particle swarm optimization to efficiently handle the associated non-convex problems [128, 129]. The wireless powered mobile edge computing in pinching-antenna system was investigated in [130].
@@ IV-C3 More Applications of Pinching Antennas
Recent advances have also demonstrated that pinching-antenna systems hold significant promise for a wide range of emerging wireless applications beyond classical communications. For example, the authors of [135] introduced a novel pinching-antenna-assisted index modulation scheme, where both conventional modulation symbols and the indices of activated pinching-antenna positions are used for information delivery. This approach enables higher spectral efficiency without increasing hardware complexity, and substantially outperforms conventional modulation systems, highlighting the versatility and potential of pinching antennas in enabling advanced modulation and coding techniques. Pinching antenna-assisted symbiotic radio was studied in [136], aiming to enhance both primary and secondary transmissions by leveraging the flexibility of pinching antennas. The authors of [137] investigated the pinching-antenna enabled indoor positioning.
@@ V Charting the Future: Open Questions and Research Directions
V-A Pinching-Antenna System Design under Random LoS/NLoS Channels
LoS blockage is a major challenge in high-frequency wireless systems, particularly in complex indoor environments, dense urban areas, and tunnel-like structures. Obstacles and user mobility can obstruct the direct propagation path between transmitters and receivers, leading to significant signal attenuation and degraded communication performance. These effects are especially pronounced in systems that rely on direct-path transmission. As pinching-antenna systems are envisioned for deployment in such environments, understanding, mitigating, and exploiting the impact of both LoS blockage and NLoS becomes essential for robust system design.